OpenStudy (anonymous):

Which answer shows this equation in standard form? 6 – 2(x – y) = 3x + 5 A.5x – 2y = 1 B.5x – 2y = –1 C.5x + 2y = 1 D.2x – 5y = 1

OpenStudy (lochana):

how did you start it? what was your fist step?

OpenStudy (anonymous):

I dont really know

OpenStudy (lochana):

fist expand brackets

OpenStudy (lochana):

\(6 - 2x +2y = 3x + 5\)

OpenStudy (anonymous):

oh ok

OpenStudy (lochana):

yea. can you do it now?

OpenStudy (lochana):

you need to arrange x and y parts left side and others right side

OpenStudy (anonymous):

idk wat to do first

OpenStudy (lochana):

sorry about delay. \(6 -2x + 2y + 2x = 3x+ 5 + 2x \\6+2y = 5x + 5 \\ 6+ 2y - 5 = 5x + 5 -5 \\ 1 + 2y = 5x \\1 +2y -2y = 5x -2y \\1 = 5x -2y \)